OpenStudy (anonymous):

How would I approach this problem? \[\sum{\frac{n^2*(-1)^n}{n!}}\] I'm trying to find the value of the series.

OpenStudy (anonymous):

series yeah!

OpenStudy (anonymous):

A hint given is that \[e^{-1} = \sum{\frac{(-1)^n}{n!}}\]

OpenStudy (anonymous):

You have to use on of the tests to figure out whether its convergent or divergent. I think ratio test would work for this.

OpenStudy (anonymous):

I know the series converges, but the problem asks to find the value of the series.

OpenStudy (anonymous):

Here's what I tried: \[\begin{align*}\sum_{n=0}^\infty \frac{n^2(-1)^n}{n!}&=\sum_{n=0}^\infty \frac{n(-1)^n}{(n-1)!}\\ &=\sum_{k=-1}^\infty \frac{(k+1)(-1)^{k+1}}{k!}&\text{substitution, }k=n-1\\ &=\sum_{k=0}^\infty \frac{(k+1)(-1)^{k+1}}{k!}\\ &=\sum_{k=0}^\infty \frac{k(-1)^{k+1}}{k!}+\sum_{k=-1}^\infty \frac{(-1)^{k+1}}{k!}\\ &=-\sum_{k=0}^\infty \frac{k(-1)^{k}}{k!}-\sum_{k=-1}^\infty \frac{(-1)^{k}}{k!}\\ &=-\color{red}{\sum_{k=0}^\infty \frac{k(-1)^{k}}{k!}}-\frac{1}{e} \end{align*}\] Checking on wolfram, you'd find that the red sum comes to \(-\dfrac{1}{e}\), but I'm assuming that's not a valid reason for the assignment. Well, continuing with this sort of decomposition, we come to some problems: \[\begin{align*} \sum\cdots &=-\sum_{k=0}^\infty \frac{k(-1)^{k}}{k!}-\frac{1}{e}\\ &=-\sum_{k=0}^\infty \frac{(-1)^{k}}{(k-1)!}-\frac{1}{e}\\ &=-\sum_{m=-1}^\infty \frac{(-1)^{m+1}}{m!}-\frac{1}{e}&\text{substituting, }m=k-1\\ &=\sum_{\color{red}{m=-1}}^\infty \frac{(-1)^{m}}{m!}-\frac{1}{e} \end{align*}\] That first \(m=-1\) term throws everything off...

OpenStudy (anonymous):

Well, as it turns out, the sum from \(m=-1\) to \(\infty\) works out the same as from \(m=0\) to \(\infty\). Hmm...

OpenStudy (anonymous):

Yeah, for some reason, \(\dfrac{(-1)^{-1}}{(-1)!}=0\)... I guess the reasoning works out.